# Bayesian Optimization of Laser-Wakefield Acceleration via Spectral Pulse Shaping

### v1

@@ Abstract
In this paper, we investigate the effect of spectral pulse shaping of the laser driver on the performance of channel-guided, laser-plasma accelerators. The study was carried out with the assistance of Bayesian optimization using particle-in-cell simulations. We used a realistic plasma profile based on a novel optical-field-ionized channel technique with ionization injection and low on-axis plasma densities to maximize the energy gain of the electron bunch trailing the laser. Spectral shaping allows us to modify the temporal profile of the laser driver while keeping the laser energy constant, affecting the acceleration and injection processes. Given the complexity and breadth of the parameter space in question, we used numerical optimization to identify high performers. In particular, we found laser profiles with additional spectral content that, when used with optimal plasma channel parameters, result in charge content an order of magnitude higher than the baseline Gaussian case while also increasing the mean energy of the electron bunch.
@@ I Introduction
Laser wakefield acceleration (LWFA) has emerged as a leading platform for generating compact, high-gradient particle acceleratorsesarey2009. In LWFA, an intense laser pulse propagating through a plasma drives a wake capable of accelerating charged particles to gigaelectronvolt (GeV) energies over tens of centimeters.akhiezer1956; tajima1979; picksley2024 This technique promises revolutionary applications in particle physics,schroeder2010; schroeder2016 medical imaging,cole2015, light sources,albert2016; barber2025 and high-energy-density science.albert2024 A recent development is the use of LWFA generated electrons as a source for muons to be used as an advanced diagnostic.terzani2025; ludwig2025
@@
Over the past two decades, LWFA has seen rapid progress in experimental performance.geddes2004; leemans2006; gonsalves2019; picksley2024 Key to these advances has been the development of tailored plasma channels that guide the laser pulse and sustain an accelerating and focusing structure.durfee1993; geddes2004; gonsalves2019 Recent innovations such as hydrodynamic optical-field-ionized (HOFI) plasma channels have enabled unprecedented control over the laser-plasma interaction, reducing diffraction and permitting longer acceleration lengths shalloo2020; miao2025. These advances have culminated in the demonstration of electron energies of 10 GeV in a single LWFA stage  picksley2024, establishing a new benchmark in the field.
@@
Despite these achievements, improving beam quality and maximizing injected charge remain major challenges. The complex, high-dimensional parameter space associated with LWFA, e.g., laser pulse shape, plasma profile, injection dynamics, etc., makes systematic optimization difficult using traditional trial-and-error approaches or grid scans. As a result, numerical optimization techniques, particularly those based on machine learning, have gained traction.
@@
Bayesian optimization (BO) with Gaussian process (GP) regression offers a powerful framework for guiding simulations or experiments toward optimal configurations with minimal evaluations.frazier2018 By building surrogate models of the objective landscape, GP-based optimization can identify trends, extrapolate performance, and efficiently navigate sparse parameter spaces. Unlike other techniques such as neural networks or random forests, GPs come with built-in uncertainty quantification, where the posterior variance tells us how uncertain the model is at any point. This approach has been successfully applied in plasma-based acceleration to optimize laser phase, experimental spectral shaping, simulations, and plasma density profiles.ziegler2021; jalas2021; loughran2023; valenta2025; mariscal2024
@@
In this work, we apply Bayesian optimization to the problem of tailoring both the laser driver profile and plasma channel to maximize the quality of the accelerated bunch. Our focus is on enhancing the bunch charge and energy using a spectrally-shaped laser pulse coupled to a tapered HOFI plasma channel. We leverage Gaussian process surrogates to explore the parameter space and identify configurations that outperform conventional laser drivers. The results demonstrate the promise of data-driven optimization for advancing the next generation of compact, tunable particle accelerators.
@@ II.1 Spectral Shaping of Short Pulse Lasers
By controlling $\phi(\omega)$, one can generate a wide variety of temporal profiles, chirped pulses, asymmetric shapes, double peaks, pulses with extended pedestals, etc. Theoretically, this is done without altering the total energy but, in reality, the spectrum can be truncated and the energy reduced. This is regularly done in experiments using a device known as an acousto-optic programmable dispersive filter (AOPDF) or DAZZLER, which is typically used to maximize compression and thereby the intensity. In this work we will only manipulate the 2nd, 3rd, and 4th order spectral coefficients since this is what is typically done in experiments, but technically we can go to arbitrary order of spectral phase.
Figure 1: Depiction of the process of spectral shaping. In (a) we can use the physical BELLA spectrum (blue), b) convolve it with a polynomial based on the Taylor expansion of the phase where we chose the coefficients of the 2nd, 3rd, and 4th order terms, and take the inverse Fourier transform to derive a new, spectrally shaped temporal profile in (c). Keeping the same polynomial phase but using idealized spectra of a Gaussian (orange) or super-Gaussian (green), we can see some variation to the final temporal profile.
@@ II.2 Spectral Shaping of Wakefield
We can examine these effects with some basic theory to build intuition. The temporal intensity profile of the laser pulse directly affects the dynamics of plasma wave excitation and electron injection. The ponderomotive force $F_{p}\propto\nabla|\mathcal{E}(t)|^{2}$ governs the wakefield amplitude and the plasma response can affect the intensity profile. Starting in the quasi-linear limit $(a_{0}^{2}\ll 1)$, the on-axis wake potential $\phi(\zeta,z)$ obeys the following relations,
@@
where $S(\zeta,z)\propto\partial_{\zeta}^{2}|a(\zeta,z)|^{2}$ is the source term, $\zeta=ct-z$, and $k_{p}=\omega_{p}/c$. Taking the Green’s function solution $\sin(k_{p}(\zeta-\zeta^{\prime})]$,esarey2009 we can express the wake potential via the causal convolution
@@
To leading order, the source term is proportional to the ponderomotive force, i.e., $F_{p}(\zeta,z)\propto\frac{\partial}{\partial\zeta}|a(\zeta,z)|^{2}.$ Letting the complex laser envelope at $z=0$ be specified by amplitude $a(\omega)$ and spectral phase $\phi(\omega)$. After propagating a distance $z$ in a plasma with dispersion $k(\omega)$, the slowly-varying envelope in the co-moving frame is $a(\zeta,z)=\mathfrak{F}^{-1}\left[a(\omega)e^{i\phi(\omega)}e^{ik(\omega)z}\right],$ where $\mathfrak{F}^{-1}$ denotes the inverse Fourier transform with respect to $\omega$ into $\zeta$. The general source entering the Green’s function integral is then
@@
We can analytically derive the effect of spectral shaping when considering a Gaussian spectral amplitude, which is not realistic, centered at $\omega_{0}$ with a bandwidth of $\sigma_{\omega}$, i.e., $a_{0}\exp(-\omega^{2}/2\sigma_{\omega}^{2})$ and quadratic spectral phase (linear chirp) $\phi(\omega)=\frac{1}{2}\phi^{\prime\prime}(\omega-\omega_{0})^{2}$. We can also expand the dispersion as $k(\omega)\approx k_{0}(\omega_{0})+k^{\prime}(\omega_{0})(\omega-\omega_{0})+\frac{1}{2}k^{\prime\prime}(\omega_{0})(\omega-\omega_{0})^{2}$, where $k^{\prime}(\omega_{0})=k_{1}=v_{g}^{-1}=\frac{dk}{d\omega}|_{\omega_{0}}$ is the inverse group velocity and $k^{\prime\prime}(\omega_{0})=\beta_{2}=\frac{d^{2}k}{d\omega^{2}}|_{\omega_{0}}$ is the group velocity dispersion coefficient.agrawal2013 Up to the carrier phase we can write the inverse Fourier transformed pulse as
@@ III Bayesian Optimization with Gaussian Processes
Optimization in laser-plasma acceleration is inherently challenging due to the high dimensionality of the parameter space, non-linear couplings, and the high cost of simulations or experimental evaluations. Traditional grid searches or manual parameter scans quickly become infeasible in such regimes. To address this, Bayesian optimization (BO) offers a powerful framework to efficiently explore complex parameter spaces with limited evaluations.frazier2018
Bayesian optimization is a surrogate-based global optimization method that probabilistically models the unknown objective function using a Gaussian process (GP). At its core, a GP defines a distribution over functions characterized by a mean function $\mu(x)$ and a covariance kernel $k(x,x^{\prime})$:
@@
Given a set of observations, the GP posterior provides both a prediction and an uncertainty estimate at the new query points. At each iteration, the GP is used to construct an acquisition function $\alpha(x)$ that balances exploration (sampling in uncertain regions) and exploitation (sampling near known high-performing points). For this work, we employ Expected Improvement (EI):
${EI}=[\mu(x)-f_{\text{best}}]\Phi(Z)+\sigma(x)\phi(Z),\ \ \ \ \ \ Z=\frac{\mu(x)-f_{\text{best}}}{\sigma(x)},$
where for a predictive distribution for $f(x)$, characterized by mean $\mu(x)$ and variance $\sigma^{2}(x)$, $f_{\text{best}}$ is the best observed value thus far and $\Phi$ and $\phi$ are the cumulative and probability distribution functions, respectively. The next evaluation point maximizes this acquisition function: $x_{\text{next}}=\arg\max_{x}\alpha(x)$. $EI$ trades off exploitation (large $\mu-f_{\text{best}}$) and exploration (large $\sigma$), naturally directing sampling towards regions that are promising or uncertain.
This approach has been successfully applied in plasma-based acceleration to optimize laser phase, spectral shaping in experiments, and plasma density profiles.jalas2021; loughran2023 In this work, we apply BO to identify optimal spectral phase profiles and plasma configurations that maximize key LWFA metrics. Each configuration is evaluated through WarpX simulations that return scalar figures of merit. The GP surrogate interpolates smoothly in low-data regimes, which is critical given the computational cost of full LWFA simulations.
In this study, we used the Scikit-Learn Bayesian optimization framework in addition to the Surrogate Modeling Toolbox.saves2024smt For deeper mathematical treatments, we refer readers to comprehensive reviews on Gaussian processes and Bayesian optimization.frazier2018; jalas2021; loughran2023; Roussel2024Bayesian
@@ IV Simulation Setup
In this study, we used the WarpX particle-in-cell (PIC) code to model laser wakefield acceleration. WarpX is a highly-parallelized and highly-optimized PIC code that can run on GPU and multi-core CPUs and includes load balancing capabilities.Fedeli2022 These simulations were run on the LLNL Lassen supercomputer and Perlmutter supercomputer at NERSC (LBNL). In addition, WarpX includes many physics and computational packages. Here we took advantage of the boosted-frame capability that allows us to relax the temporal resolution requirements of the simulation given that distances are scaled by the relativistic Lorentz factor $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The computational gain comes from the fact that in the boosted frame the laser is redshifted. The scale imbalance that results in long simulations in the lab frame is reduced (the plasma is shorter and the laser wavelength is longer). In addition, we use the LASY laser package, which allows laser field and phase profiles to be inputted into the WarpX particle-in-cell code, as well as forward and backward propagate the laser pulse so that it can be properly injected into the WarpX simulation at the appropriate focus point.
@@
In this study, our simulations have the following numerical setup. The simulation bounds are $256,0]\times[0,128]\mu m$ with a grid spacing of $10,240]\times[128]$ for the longitudinal and horizontal coordinates, which means that the spatial resolution is $0.025\ \mu m$ and $1\ \mu m$, respectively. The laser wavelength, a free variable in this study, is approximately 800 nm, meaning that we have 32 grid cells per laser wavelength longitudinally, which was found to be sufficient after extensive convergence testing and calibration. We use the boosted frame moving window package with a gamma boost factor of $\gamma=10$; further enhancement leads to numerical artefacts forming that are considered to be nonphysical. Damped boundary conditions are used on the longitudinal boundaries of the simulation box, which applies a damping factor to the electric and magnetic fields in the outer half of the guard cells, using a sine-squared profile. This is necessary since we used the Pseudo Spectral Analytical Time Domain (PSATD) module to solve our system in RZ coordinates. We use 3 azimuthal modes and a momentum-conserving algorithm for field gathering. Particle count is varied across the length of the simulation, with higher counts in the initial, injection region of the plasma column, $(\text{ppc}_{r},\text{ppc}_{\theta},\text{ppc}_{z})=(3,14,3)$, which is then relaxed after the laser passes through the dopant region and there is little injection still occurring, i.e., $(2,8,2)$. This was necessary to have convergent results for the electron bunch primarily via ionization injection but also to reduce the overall computational cost of the simulation.
@@
The physical setup of our target problem is as follows. The laser driver is described by disjoint temporal and transverse functions, $a(r,\tau)=a_{r}(r)a_{\tau}(t,\text{GDD,TOD,FOD})$. Transversely it is described a Jinc function, i.e., $a_{r}(r)=\frac{2J_{1}(r/r_{w})}{(r/r_{w})}$, which describes an initially far-field top-hat laser profile as it reaches focus, as seen in Fig. 2.a), where we compare it to a Gaussian transverse profile and a transverse slice of the plasma column. Here $r_{w}=r_{0}/2.66$, where $r_{0}=52\ \mu m$ is the equivalent transverse pulse radius with respect to the Gaussian full-width at half maximum. Using LASY we can easily back-propagate a Jinc type laser so as to properly focus it into the target plasma. The temporal profile is described by 3 spectral components, i.e., GDD, TOD, and FOD, as previously described. The laser has a constant 30 J energy for all pulse shapes, although in reality spectral shaping can effectively truncate parts of the spectrum so that the laser’s intensity is less than what was theoretically predicted.
@@
The plasma column through which the laser propagates is more complex than normally used. Here we used a custom plasma profile based on FLASH radiation hydrodynamics simulations that mimic the type of profile we expect to see given the HOFI plasma channel used in recent experiments.diaw2022; cook2020; cook2024; cook2025a HOFI uses a preceding heater beam to generate the effectively quasi-parabolic profile to guide the drive laser over long distances. An example plasma profile can be seen in Fig. 2.a). In addition, we have included longitudinal structure to the plasma channel to help optimize LWFA outputs. We typically have a plasma channel length of 60 cm in our simulations, although significant dephasing and depletion can begin as early as around 30 cm. The first few centimeters of the plasma is composed of a nitrogen-doped hydrogen gas, where the nitrogen is a few percent of the channel density. The hydrogen is fully ionized and the nitrogen is partially ionized to $N^{5+}$. As the laser propagates through the dopant layer it liberates $N^{6+}$ and $N^{7+}$ electrons, which constitute the primary contributions to the acceleration electron bunch, few background electrons feed into the final electron bunch. WarpX uses the Ammosov-Delone-Krainov (ADK) tunnel-ionization model. In addition, we allow the plasma profile to have a density taper, where the density may increase linearly along the propagation axis $z$. A visualization of such a profile can be found in Fig. 2.b).

### v2

@@ Abstract
In this paper, we investigate the effect of spectral pulse shaping of the laser driver on the performance of channel-guided, laser-plasma accelerators. The study was carried out with the assistance of Bayesian optimization using particle-in-cell simulations. We used a realistic plasma profile based on a novel optical-field-ionized channel technique with ionization injection and low, on-axis plasma densities to maximize the energy gain of the electron bunch trailing the laser. Spectral shaping allows us to modify the temporal profile of the laser driver while keeping the laser energy constant, affecting the acceleration and injection processes. In addition we consider how modifying the plasma channel parameters may affect the target outputs. Given the complexity and breadth of the parameter space in question, we used numerical optimization to identify high performers. In particular, we found laser profiles with additional spectral content that, when used with optimal plasma channel parameters, result in charge content an order of magnitude higher than the baseline Gaussian case while also increasing the mean energy of the electron bunch.
@@ I Introduction
Laser wakefield acceleration (LWFA) has emerged as a leading platform for generating compact, high-gradient particle acceleratorsakhiezer1956; tajima1979; esarey2009. In LWFA, an intense laser pulse propagating through a plasma drives a wake capable of accelerating charged particles to gigaelectronvolt (GeV) energies over tens of centimeters.picksley2024; rockafellow2025 This technique promises revolutionary applications in particle physics,schroeder2010; schroeder2016 medical imaging,cole2015, light sources,albert2016; barber2025 and high-energy-density science.albert2024 A recent development is the use of LWFA generated electrons as a source for muons to be used as an advanced diagnostic.terzani2025; ludwig2025
@@
Over the past two decades, LWFA has seen rapid progress in experimental performancegeddes2004; leemans2006; gonsalves2019; miao2022 modeling capability,lehe2015; vay2016; shapoval2024 and the underlying theory.lu2006; tzoufras2008; esarey2009 Key to these advances has been the development of tailored plasma channels that guide the laser pulse and sustain an accelerating and focusing structure.durfee1993; geddes2004; gonsalves2019 Recent innovations such as hydrodynamic optical-field-ionized (HOFI) plasma channels have enabled unprecedented control over the laser-plasma interaction, reducing diffraction and permitting longer acceleration lengths shalloo2020; miao2020. These advances have recently culminated in the demonstration of electron energies of 10 GeV in a single LWFA stage  picksley2024; rockafellow2025; aniculaesei2023, establishing a new benchmark in the field.
@@
Despite these achievements, improving beam quality and maximizing injected charge remain major challenges. The complex, high-dimensional parameter space associated with LWFA, e.g., laser pulse shape, plasma profile, injection dynamics, etc., makes systematic optimization difficult using traditional trial-and-error approaches or grid scans. As a result, numerical optimization techniques, particularly those based on machine learning, have gained traction. Controlling the spectral phase of the laser pulse for short-pulse laser plasma interactions in particular, whether to maximize the compression or tune the pulse shape, is challenging to do manually and has benefited from various numerical optimization techniques already.streeter2018; dann2019; shalloo2020; mariscal2024
@@
Bayesian optimization (BO) with Gaussian process (GP) regression offers a powerful framework for guiding simulations or experiments toward optimal configurations with minimal evaluations.frazier2018 The algorithm iteratively: (1) fits a Gaussian process to all available observations, (2) uses this GP to define an acquisition function that quantifies the expected utility of sampling at each point, and (3) evaluates the objective at the point maximizing this acquisition function. By building surrogate models of the objective landscape, GP-based optimization can identify trends, extrapolate performance, and efficiently navigate sparse parameter spaces. Unlike other techniques such as neural networks or random forests, GPs come with built-in uncertainty quantification, where the posterior variance tells us how uncertain the model is at any point. This approach has been successfully applied in plasma-based acceleration to optimize laser phase, experimental spectral shaping,ziegler2021; loughran2023 simulations,valenta2025 injection mechanisms,backhouse2025 plasma density profiles,jalas2021 and multi-objective optimization of LWFAs.irshad2023
@@
In this work, we apply Bayesian optimization to the problem of tailoring both the laser driver profile and plasma channel to maximize the quality of the accelerated bunch. Our focus is on enhancing the bunch charge and energy while minimizing energy spread using a spectrally-shaped laser pulse coupled to a tapered HOFI plasma channel. We leverage Gaussian process surrogates to explore the parameter space and identify configurations that outperform conventional laser drivers. The results demonstrate the promise of data-driven optimization for advancing the next generation of compact, tunable particle accelerators.
@@ II.1 Spectral Shaping of Short Pulse Lasers
By controlling $\phi(\omega)$, one can generate a wide variety of temporal profiles, chirped pulses, asymmetric shapes, double peaks, pulses with extended pedestals, etc. This is regularly done in experiments using a device known as an acousto-optic programmable dispersive filter (AOPDF) or DAZZLER™, which is typically used to maximize compression and thereby the intensity. Theoretically, this is done without altering the total energy but, in practice, the spectrum can be truncated and the energy reduced. In this work we will only manipulate the 2nd, 3rd, and 4th order spectral coefficients since this is what is typically done in experiments, but technically we can go to arbitrary order of spectral phase.
Figure 1: Depiction of the process of spectral shaping. In (a) we can use the physical BELLA spectrum (blue), convolve it with b) a polynomial based on the Taylor expansion of the phase, where we chose the coefficients of the 2nd, 3rd, and 4th order terms, and take the inverse Fourier transform to derive a new, spectrally shaped temporal profile in (c). Keeping the same polynomial phase but using idealized spectra of a Gaussian (orange) or super-Gaussian (green), we can see some variation to the final temporal profile. We also consider a skewed spectrum (red), which can be seen in experiments.
@@ II.2 Spectral Shaping of Wakefield
We can examine these effects with some basic theory to build intuition. The temporal intensity profile of the laser pulse directly affects the dynamics of plasma wave excitation and electron injection. The ponderomotive force $F_{p}\propto\nabla|\mathcal{E}(t)|^{2}$ governs the wakefield amplitude and the plasma response can affect the intensity profile. Starting in the quasi-linear limit $(a_{0}^{2}\ll 1)$, the on-axis wake potential $\phi(\zeta,z)$ obeys the following relation,
@@
where $S(\zeta,z)\propto\partial_{\zeta}^{2}|a(\zeta,z)|^{2}$ is the source term, $\zeta=ct-z$, and $k_{p}=\omega_{p}/c$. Taking the Green’s function solution $\sin[k_{p}(\zeta-\zeta^{\prime})]$,esarey2009 we can express the wake potential via the causal convolution
@@
To leading order, the source term is proportional to the ponderomotive force, i.e., $F_{p}(\zeta,z)\propto\frac{\partial}{\partial\zeta}|a(\zeta,z)|^{2}.$ We let the complex laser envelope at $z=0$ be specified by amplitude $a(\omega)$ and spectral phase $\phi(\omega)$. After propagating a distance $z$ in a plasma with dispersion $k(\omega)$, the slowly-varying envelope in the co-moving frame is $a(\zeta,z)=\mathfrak{F}^{-1}\left[a(\omega)e^{i\phi(\omega)}e^{ik(\omega)z}\right],$ where $\mathfrak{F}^{-1}$ denotes the inverse Fourier transform with respect to $\omega$ onto $\zeta$. The general source entering the Green’s function integral is then
@@
We can analytically derive the effect of spectral shaping when considering a Gaussian spectral amplitude centered at $\omega_{0}$ with a bandwidth of $\sigma_{\omega}$, i.e., $a_{0}\exp(-\omega^{2}/2\sigma_{\omega}^{2})$ and quadratic spectral phase (linear chirp) $\phi(\omega)=\frac{1}{2}\phi^{\prime\prime}(\omega-\omega_{0})^{2}$. We can also expand the dispersion as $k(\omega)\approx k_{0}(\omega_{0})+k^{\prime}(\omega_{0})(\omega-\omega_{0})+\frac{1}{2}k^{\prime\prime}(\omega_{0})(\omega-\omega_{0})^{2}$, where $k^{\prime}(\omega_{0})=k_{1}=v_{g}^{-1}=\frac{dk}{d\omega}|_{\omega_{0}}$ is the inverse group velocity and $k^{\prime\prime}(\omega_{0})=\beta_{2}=\frac{d^{2}k}{d\omega^{2}}|_{\omega_{0}}$ is the group velocity dispersion coefficient.agrawal2013 Up to the carrier phase we can write the inverse Fourier transformed pulse as
@@ III Bayesian Optimization with Gaussian Processes
Optimization in laser-plasma acceleration is inherently challenging due to the high dimensionality of the parameter space, non-linear effects, and the high cost of simulations and experimental evaluations. Traditional grid searches or manual parameter scans quickly become infeasible in such regimes. To address this, Bayesian optimization (BO) offers a powerful framework to efficiently explore complex parameter spaces with limited evaluations.frazier2018
BO is a surrogate-based global optimization method that probabilistically models an unknown objective function using a Gaussian process (GP). GP regression produces a posterior probability distribution where each $f(x)$ is normally distributed with mean $\mu(x)$ and variance $\sigma^{2}(x)$, updated after $n$ observations. The mean can be interpreted as a point estimate of $f(x)$, while the variance quantifies our uncertainty. At its core, a GP defines a distribution over functions characterized by a mean function $\mu(x)$ and a covariance kernel $k(x,x^{\prime})$:
@@
Given a set of observations, the GP posterior provides both a prediction and an uncertainty estimate at the new query points. At each iteration, the GP is used to construct an acquisition function $\alpha(x)$ that balances exploration (sampling in more uncertain regions) and exploitation (sampling near known high-performing points) efficiently locate the global optimum. For this work, we employ the standard acquisition function of Expected Improvement (EI):
${EI(x)}=[\mu(x)-f_{\text{best}}]\Phi(Z)+\sigma(x)\phi(Z),\ \ \ \ \ \ Z=\frac{\mu(x)-f_{\text{best}}}{\sigma(x)},$
again, where for a predictive distribution for $f(x)$, characterized by mean $\mu(x)$ and variance $\sigma^{2}(x)$, $f_{\text{best}}$ is the best observed value thus far and $\Phi$ and $\phi$ are the cumulative and probability distribution functions, respectively. The next evaluation point maximizes this acquisition function: $x_{\text{next}}=\arg\max_{x}\alpha(x)$. $EI$ trades off exploitation (large $\mu-f_{\text{best}}$) and exploration (large $\sigma$), naturally directing sampling towards regions that are promising or uncertain. While GP regression scales as $O(n^{3})$ with the number of observations, this is acceptable for expensive black-box functions where $n$ typically remains below a few hundred evaluations
In this work, we apply BO to identify optimal spectral phase profiles and plasma configurations that maximize key LWFA metrics. Each configuration is evaluated through WarpX simulations that return scalar figures of merit. The GP surrogate interpolates smoothly in low-data regimes, which is critical given the computational cost of full LWFA simulations. Here we used the Scikit-Learn Bayesian optimization framework in addition to the Surrogate Modeling Toolbox.saves2024smt For deeper mathematical treatments, we refer readers to comprehensive reviews on Gaussian processes and Bayesian optimization.frazier2018; Roussel2024Bayesian
@@ IV Simulation Setup
In this study, we used the WarpX particle-in-cell (PIC) code to model laser wakefield acceleration. WarpX is a highly parallelized and optimized PIC code that can run on GPU and multi-core CPUs and includes load balancing capabilities.Fedeli2022 These simulations were run on the LLNL Lassen supercomputer and Perlmutter supercomputer at NERSC (LBNL). In addition, WarpX includes many physics and computational packages. Here we took advantage of the boosted-frame capability that allows us to relax the temporal resolution requirements of the simulation given that distances are scaled by the relativistic Lorentz factor $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The computational gain comes from the fact that in the boosted frame the laser is redshifted. The scale imbalance that results in long simulations in the lab frame is reduced (the plasma is shorter and the laser wavelength is longer). In addition, we use the LASY laser package, which allows laser field and phase profiles to be inputted into the WarpX particle-in-cell code, as well as forward and backward propagate the laser pulse so that it can be properly injected into the WarpX simulation at the appropriate focus point.
@@
In this study, our simulations have the following numerical setup. The simulation bounds are $256,0]\times[0,128]\mu m$ with a grid discretization of $10,240]\times[128]$ for the longitudinal and horizontal coordinates, which means spatial resolutions of $0.025\ \mu m$ and $1\ \mu m$, respectively. The laser wavelength, a free variable in this study, is approximately 800 nm, meaning that we have 32 grid cells per laser wavelength longitudinally, which was found to be sufficient after extensive convergence testing and calibration. We use the boosted frame moving window package with a gamma boost factor of $\gamma=10$; further enhancement leads to numerical artefacts forming that are considered to be nonphysical. Damped boundary conditions are used on the longitudinal boundaries of the simulation box, which applies a damping factor to the electric and magnetic fields in the outer half of the guard cells, using a sine-squared profile. This is necessary since we used the Pseudo Spectral Analytical Time Domain (PSATD) module to solve our system in RZ coordinates. We use 3 azimuthal modes and a momentum-conserving algorithm for field gathering. Particle count is varied across the length of the simulation, with higher counts in the initial, injection region of the plasma column, $(\text{ppc}_{r},\text{ppc}_{\theta},\text{ppc}_{z})=(3,14,3)$, which is then relaxed after the laser passes through the dopant region, after which there is little injection still occurring, i.e., $(2,8,2)$. This was necessary to have convergent results for the electron bunch primarily via ionization injection but also to reduce the overall computational cost of the simulation.
@@
The physical setup of our target problem is as follows. The laser driver is described by disjoint temporal and transverse functions, $a(r,\tau)=a_{r}(r)a_{\tau}(t,\text{GDD,TOD,FOD})$. Transversely it is described a Jinc function, i.e., $a_{r}(r)=\frac{2J_{1}(r/r_{w})}{(r/r_{w})}$, which describes an initially far-field top-hat laser profile as it reaches focus, as seen in Fig. 2.a), where we compare it to a Gaussian transverse profile and a transverse slice of the plasma column. Here $r_{w}=r_{0}/2.66$, where $r_{0}=52\ \mu m$ is the equivalent transverse pulse radius with respect to the Gaussian full-width at half maximum. Using LASY we can easily back-propagate a Jinc type laser so as to properly focus it into the target plasma. The temporal profile is described by 3 spectral components, i.e., GDD, TOD, and FOD, as previously described. The laser has a constant 30 J energy for all pulse shapes, although in experiments spectral shaping can effectively truncate parts of the spectrum so that the laser’s intensity is less than what was theoretically predicted.
@@
The plasma column through which the laser propagates is more complex than normally used. Here we used a custom plasma profile based on FLASH radiation hydrodynamics simulations that mimic the type of profile we expect to see given the HOFI plasma channel used in recent experiments.diaw2022; cook2020; cook2024; cook2025a HOFI uses a preceding heater beam to generate the plasma profile to guide the drive laser over long distances. An example plasma profile can be seen in Fig. 2.a). In addition, we have included longitudinal structure to the plasma channel to help optimize LWFA outputsdephasing. We typically have a plasma channel length of 60 cm in our simulations, although significant dephasing and depletion can begin as early as around 30 cm. The first few centimeters of the plasma is composed of a nitrogen-doped hydrogen gas, where the nitrogen is a few percent of the channel density. The hydrogen is fully ionized and the nitrogen is partially ionized to $N^{5+}$. As the laser propagates through the dopant layer it liberates $N^{6+}$ and $N^{7+}$ electrons, which constitute the primary contributions to the acceleration electron bunch, few background electrons feed into the final electron bunch. WarpX uses the Ammosov-Delone-Krainov (ADK) tunnel-ionization model. In addition, we allow for the plasma profile to have a density taper, where the density may increase linearly along the propagation axis $z$, which can be used to compensate for dephasing.rittershofer2010; djordjevic2019 A visualization of such a profile can be found in Fig. 2.b).